Johnson (1966) measured bills of dusky flycatchers and found that bill lengths of the
males had a mean of 8.15 ± 0.021 mm (mean ± SE) and a coefficient of variation of
4.67 %. On the basis of this information, infer how many specimens must have been
used in the sample.
1
Expert's answer
2020-04-19T08:10:49-0400

Coefficient of Variation "=\frac{s} {\bar x} *100"

4.67="\frac {s} {8.15}*100"

s=0.3806

SE="\frac {s} {\sqrt n}"

0.021="\frac {0.3806}{\sqrt n}"

n="(\frac {0.3806}{0.021})^2"

The sample consists of n=329 speciments.